### v1

# SenseFlow: Scaling Distribution Matching for Flow-based Text-to-Image Distillation

Abstract
The Distribution Matching Distillation (DMD) has been successfully applied to text-to-image diffusion models such as Stable Diffusion (SD) 1.5. However, vanilla DMD suffers from convergence difficulties on large-scale flow-based text-to-image models, such as SD 3.5 and FLUX. In this paper, we first analyze the issues when applying vanilla DMD on large-scale models. Then, to overcome the scalability challenge, we propose implicit distribution alignment (IDA) to regularize the distance between the generator and fake distribution. Furthermore, we propose intra-segment guidance (ISG) to relocate the timestep importance distribution from the teacher model. With IDA alone, DMD converges for SD 3.5; employing both IDA and ISG, DMD converges for SD 3.5 and FLUX.1 dev. Along with other improvements such as scaled up discriminator models, our final model, dubbed SenseFlow, achieves superior performance in distillation for both diffusion based text-to-image models such as SDXL, and flow-matching models such as SD 3.5 Large and FLUX. The source code will be avaliable at https://github.com/XingtongGe/SenseFlow.
Figure 1: 1024×1024 samples produced by our 4-step generator distilled from FLUX.1-dev.
1 Introduction
Significant advancements have been made on diffusion models ho2020denoising ; rombach2022high ; podell2023sdxl ; esser2024scaling ; flux2024 for text-to-image generation over recent years. However, these models typically require multiple denoising steps to generate high-quality images. As models continue to scale up in terms of the parameter size, the computational cost and inference time for image generation increase substantially, making the process slower and more resource-intensive. To address this issue, various diffusion distillation methods have been developed to distill a diffusion model into a few-step generator, including consistency models song2023consistency ; luo2023latent ; wang2024phased , progressive distillation salimans2022progressive ; ren2024hyper , adversarial distillation Wang2023DiffusionGAN ; sauer2024adversarial ; chadebec2025flash , and score distillation yin2024one ; yin2024improved ; Luo2023diffinstruct . Currently, the Distribution Matching Distillation (DMD) series yin2024improved have demonstrated superior results in distilling standard diffusion models such as SD 1.5 rombach2022high and SDXL podell2023sdxl .
However, few of these methods have successfully demonstrated effective distillation performance across a broader range of models, particularly in flow-based diffusion models with larger parameter sizes, such as SD3.5 Large (8B) esser2024scaling and FLUX.1 dev (12B) flux2024 . As models increase in architecture complexity, parameter size, and training complexity, it becomes significantly more challenging to distill these models into efficient few-step generators (e.g., a 4-step generator).
In this paper, we introduce SenseFlow, which selects the framework of DMD2 yin2024improved as a touchstone, and scales it up for larger flow-based text-to-image models, including SD3.5 Large and FLUX.1 dev. Specifically, vanilla DMD2 has difficulty in converging and faces significant training instability on large models, even with the time-consuming two time-scale update rule (TTUR) ttur applied. To address this challenge, we propose implicit distribution alignment (IDA) to regularize the distance between the generator and the fake distribution network, which makes the training of fake distribution network faster and easier. This further allows us to make the generator converge more stably.
Further, DMD2 and most existing diffusion distillation methods still use uniformly sampled timesteps for training and inference. However, due to the complex strategies employed during training of teacher diffusion models, different timesteps exert varying denoising effects throughout the entire process, which is also discussed in RayFlow shao2025rayflow . To avoid the inefficiency of naive timestep sampling strategy in distillation, we propose to relocate the teacher’s timestep-wise denoising importance into a small set of selected coarse timesteps. For each coarse timestep $\tau_{i}$, we construct an intra-segment guidance (ISG) by sampling an intermediate timestep $t_{mid}\in(\tau_{i-1},\tau_{i})$, and construct a guidance trajectory: the teacher denoises from $\tau_{i}$ to $t_{mid}$, then the generator continues from $t_{mid}$ to $\tau_{i-1}$. We then guide the generator to align its direct prediction from $\tau_{i}$ to $\tau_{i-1}$ with this trajectory. This guidance mechanism effectively aggregates the teacher’s fine-grained behavior within each segment, improving the generator’s ability to approximate complex transitions across fixed sparse timesteps.
For further enhancement, we incorporate a more general and powerful discriminator built upon vision foundation models (e.g., DINOv2 oquab2023dinov2 , CLIP radford2021learning ), which operates in the image domain and can provide stronger semantic guidance. The use of pretrained vision backbones introduces rich semantic priors, enabling the discriminator to better capture image-level quality and fine-grained structures. By aggregating timestep-aware adversarial signals, this design yields stable and efficient training with superior visual qualities.
To summarize, we dive into the distribution matching distillation (DMD) and scale it up for a wide range of large-size flow-based text-to-image models. Our contributions are as follows:
  • We discover that vanilla DMD2 suffers from the convergence issue on large-scale text-to-image models, even with TTUR introduced. To tackle this challenge, we propose implicit distribution alignment to regularize the distance between the generator and fake distribution.
  • To mitigate the problem of suboptimal sampling in DMD2, we propose intra-segment guidance to relocate the teacher’s timestep-wise denoising importance, improving the generator’s ability to approximate complex transitions across sparse timesteps.
  • By incorporating a more powerful discriminator built upon vision foundation models with timestep-aware adversarial signals, we achieve stable training with superior performance.
  • Experimental results show that our final model, dubbed SenseFlow, achieves superior performance in distilling large-scale flow-matching models ( e.g., SD 3.5, FLUX.1 dev) and diffusion-based models (e.g., SDXL). Our SD 3.5 Based-SenseFlow achieves state-of-the-art 4-step generation performance among all open-source models evaluated in our study.
2 Preliminaries
2.1 Diffusion Models
Diffusion models are a family of generative models, with the forward process perturbing the data $X_{0}\sim p(X_{0})$ to Gaussian noise $p(X_{T})=\mathcal{N}(0,I)$ with a series distributions $p(X_{t})$ defined by a forward stochastic differential equation (SDE):
$dX_{t}=f(X_{t},t)dt+g(t)dB_{t},t\in[0,T]$ (1)
where $f(X_{t},t)$ is drifting parameter, $g(t)$ is diffusion parameter and $B_{t}$ is standard Brownian motion. The diffusion model learns the score function $s(X_{t},t)=\nabla_{X_{t}}\log p(X_{t})$ using neural network. And the sampling of diffusion process is to solve the probability flow ordinary differential equation:
$dX_{t}=(f(X_{t},t)-\frac{1}{2}g(t)^{2}s(X_{t},t))dt,X_{T}\sim\mathcal{N}(0,I).$ (2)
The two widely adopted diffusion models in text-to-image, namely denoising diffusion probabilistic model (DDPM) and flow matching optimal transport (FM-OT), fit in above framework by setting $f(X_{t},t)=-\frac{1}{2}\beta_{t}X_{t},g(t)=\sqrt{\beta_{t}}$ and $f(X_{t},t)=-\frac{1}{1-t}X_{t},\frac{1}{2}g(t)^{2}=\frac{t}{1-t}$ respectively, where $\beta_{t}$ is hyper-parameter of DDPM. The forward SDE of DDPM and FM-OT can be directly solved:
[…]
However, the backward equation in Eq. 2 is intractable as $s(X_{t},t)$ is neural network. Usually we need time-consuming multi-step solvers. In this paper, we focus on distilling the solution of backward equations into another neural network.
2.2 Distribution Matching Distillation
From now on we assume a pre-trained diffusion model is available, with learned score function $s_{r}(X_{t},t)$ and distribution $p_{r}(X_{t})$. The Distribution Matching Distillation (DMD) yin2024one ; yin2024improved distills the diffusion model by a technique named score distillation poole2022dreamfusion . More specifically, DMD learns the generator distribution $p_{g}(X_{t})$ to match the diffusion distribution $p_{r}(X_{t})$:
$\min_{p_{g}}D_{KL}(p_{g}(X_{t})||p_{r}(X_{t}))=\mathbb{E}_{t\sim[0,T],p_{g}}[\log p_{g}(X_{t})-\log p_{r}(X_{t})].$ (5)
Directly distillation from above target produces suboptimal results. Therefore, DMD introduces an intermediate fake distribution $p_{f}(X_{t},t)$, and optimizes the generator distribution $p_{g}$ and fake distribution $p_{f}$ in an interleaved way:
[…]
In practice, the fake distribution is parameterized as the score function $s_{\phi}(X_{t},t)=\nabla\log p_{f}(X_{t})$. On the other hand, the generator is parameterized with a clean image generating network $G_{\theta}(\epsilon),\epsilon\sim\mathcal{N}(0,I)$ and forward diffusion process $q(X_{t}|X_{0})$, such that $p_{g}(X_{t})=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}[q(X_{t}|G_{\theta}(\epsilon))]$. To this end, the DMD updates are achieved by gradient descent and score matching vincent2011connection :
[…]
3 Method: Scaling Distribution Matching for General Distillation
Figure 2: Left: The generator $\mathcal{G}$ receives a text prompt and $x_{\tau_{i}}$ to produce one-step output $x_{g}$, which is diffused to $x_{t}$ and processed by $s_{\phi}$ and $s_{r}$ for computing DMD gradient. ISG guides $\mathcal{G}$ using an intermediate point $x_{t_{mid}}$, and IDA aligns $\mathcal{G}$ with $s_{\phi}$ after generator update. Right: The discriminator extracts semantic features from generated and real images using CLIP and DINOv2, which are processed by head blocks ${h_{\theta_{i}}}$ to predict real/fake logits for adversarial training. Trainable modules are shown in pink, while frozen (pretrained) ones are shown in grey.
3.1 Bottlenecks in Vanilla DMD series: Stability, sampling, and naive discriminator
While Distribution Matching Distillation (DMD) has shown promising results in aligning generative distributions, its vanilla formulation exhibits several fundamental limitations when applied to large-scale models. First, scalability remains a challenge—the two time-scale update rule (TTUR), effective in SD 1.5 (0.8B) and SDXL (2.6B), fails to converge stably when scaled to larger models such as SD 3.5 Large (8B) or FLUX (12B). Second, sampling efficiency is limited as the generator does not incorporate the varying importance of timesteps in the denoising trajectory, which slows convergence and reduces expressiveness. Third, the discriminator lacks generality, with a relatively naive design that struggles to adapt across diverse model scales and architectures. These issues motivate us to propose architectural and algorithmic improvements in this work.
3.2 Implicit Distribution Alignment via Generator-Fake Distribution Fusion
In Distribution Matching Distillation (DMD), a critical challenge lies in stabilizing the fake distribution model to accurately track the generator distribution $p_{g}$, especially when working with modern large-scale diffusion backbones such as SD3.5 or FLUX. As model capacity increases and the training strategies of teacher models vary across architectures, ensuring a well-trained fake distribution model becomes increasingly difficult. For example, many models diffusiondpo ; playgroundv3 ; seedream2 use complex post-training strategies to improve the performance of the model in specific directions, such as text rendering or aesthetic quality, which may introduce non-uniform sampling trajectories, making the standard diffusion loss less effective for supervising fake distribution model training.
Figure 3: “Training Hours-FID” curves on COCO-5K dataset. IDA improves training stability across TTUR ratios.
To address this issue, DMD2 used the two time-scale update rule (TTUR), which increases the update frequency of the fake distribution model relative to the generator. However, TTUR becomes increasingly expensive and brittle as the model size scales up. Results in Fig. 3 also indicate that sometimes even a high ratio of 20:1 still cannot stabilize the training.
On the other hand, although the generator and fake distribution network are optimized via different objectives, their long-term goals are highly aligned: both aim to model a distribution $p_{g}(X_{t})$ that closely approximates the real data distribution $p_{r}(X_{t})$. In practice, they are initialized from the same pretrained teacher and both define the generator-induced distribution $p_{g}(X_{t})$. The key difference is that, generator is guided by an explicit, fixed teacher score $s_{r}(X_{t},t)$ through the variational gradient in Eq. 7 and thus evolves in a clear direction toward $p_{r}$. In contrast, the fake distribution network is trained to regress toward the score of the generator-induced distribution via $\mathcal{L}_{f}$ in Eq. 7, where the target $\nabla\log p_{g}(X_{t})$ is approximated through the generator’s outputs. In early training, this target is a rapidly moving and highly unreliable signal—making the fake distribution network prone to underfitting, drift, or misaligned gradients, especially when the model size is relatively large.
We address this challenge by introducing Implicit Distribution Alignment (IDA), a simple yet effective stabilization mechanism. Specifically, after each generator update, we partially align the fake distribution parameters toward the generator:
$\phi\leftarrow\lambda\cdot\phi+(1-\lambda)\cdot\theta.$ (8)
Intuitively, this allows us to propagate the teacher’s stable supervision—received by the generator—into the fake distribution model indirectly. Since both networks share initialization and long-term alignment, IDA can implicitly regularize the distributional distance between the fake distribution model and generator, preventing it from being misled by the drift moving targets during early training.
In practice, this strategy ensures that the fake distribution remains closely aligned with the generator’s distributional trajectory, especially early in training when score updates are unstable. We observe that combining IDA with even a relatively small TTUR (e.g., 5:1) leads to significantly more stable convergence. An example of this effect is shown in Fig. 3, where we compare FID curves under different TTUR settings with and without IDA. As the figure illustrates, IDA consistently reduces FID variance and improves overall performance. We leave a detailed analysis to the ablation study section.
3.3 Generator Turn: Relocate the Timestep Importance Distribution
Figure 4: Left: The normalized reconstruction errors over timesteps in $[0,1]$. Right: An illustration of the Intra-Segment Guidance.
On the other hand, the distillation performance of vanilla DMD2 is fundamentally limited by fixed timestep supervision. In vanilla DMD2 setups, the generator is only trained at a small set of pre-defined timesteps (e.g., $\tau\in\{249,499,749,999\}$). However, this fixed design introduces two major issues: first, the generator receives no training signal from the rest of the trajectory, which leads to poor generalization for the full trajectory; second, the effectiveness of each supervised timestep is highly sensitive to where it lies along the trajectory—neighboring timesteps can exhibit drastically different predictive errors. To better illustrate the local reliability of different timesteps in the diffusion trajectory, we visualize the normalized one-step reconstruction loss $\xi(t)$ over 1000 uniformly spaced timesteps in $[0,1]$:
$\xi(t):=\mathbb{E}_{x_{0},\epsilon\sim\mathcal{N}(0,I)}\left[\|\hat{x}_{0}(x_{t},t)-x_{0}\|^{2}\right],$ (9)
where $x_{0}$ is generated by the teacher model (SD 3.5 or FLUX.1 dev) and $x_{t}$ is obtained via diffusion forward process in Eq. 4 using $x_{0}$ and $\epsilon$. The results are shown in Fig. 4 Left. We observe that as $t$ increases, the denoising error $\xi(t)$ does not grow monotonically, but instead exhibits noticeable local oscillations—particularly in the interval $t\in[0.8,1.0]$. This suggests that even adjacent timesteps within the same region may differ significantly in their denoising accuracy, implying that their relative “importance” to the overall denoising process is not uniform. Consequently, selecting supervision points without considering their local reliability may inadvertently anchor the generator to suboptimal points, degrading sample quality and training stability.
To mitigate this issue, we propose to relocate the teacher’s denoising importance into a small set of selected coarse timesteps. For each coarse timestep $\tau_{i}$, we construct an intra-segment guidance by randomly sampling an intermediate timestep $t_{1}\in(\tau_{i-1},\tau_{i})$. As shown in Fig. 4 Right, the teacher model generates $x_{t_{1}}$ by denoising from $\tau_{i}$ to $t_{1}$. Then, the generator continues the denoising process from $t_{1}$ to $\tau_{i-1}$, yielding the guidance target $x_{\text{tar}}$. Meanwhile, the generator also produces $x_{\tau_{i-1}}$ directly from $\tau_{i}$ to $\tau_{i-1}$. We then apply an $\mathcal{L}_{2}$ loss between $x_{g}$ and $x_{\text{tar}}$, where gradients are only propagated through the generator path:
$\mathcal{L}_{\text{ISG}}^{(i)}=\mathbb{E}_{\epsilon,t_{1}}\left[\left\|x_{g}-\mathrm{stop\_grad}(x_{\text{tar}})\right\|_{2}^{2}\right].$ (10)
This enables each anchor point to better absorb the denoising knowledge of its surrounding segment, thereby serving as a more representative proxy for its local denoising behavior.
3.4 Bonus: General and Powerful Discriminator built upon Vision Foundation Models
As shown in Fig. 2, the discriminator $D$ is designed by integrating a fixed pre-trained Vision Foundation Model (VFM) backbone, $f_{\text{VFM}}$, with learnable discriminator heads, $h$. Given an input image $x$, the VFM backbone extracts multi-level semantic features $z=f_{\text{VFM}}(x)$, which are subsequently processed by the discriminator heads to predict the realism of $x$. Additionally, the discriminator incorporates CLIP-encoded features $c=f_{\text{CLIP}}(\text{text})$ and reference features $r=f_{\text{VFM}}(x)$ from real images to additionally impregnate text-image alignment information. This process is expressed as:
$D(x)=h(f_{\text{VFM}}(x),c,r).$ (11)
These features enhance the discriminator’s capacity to evaluate both the realism and semantic consistency of the input images. The discriminator is trained using the hinge loss, defined as:
$\mathcal{L}_{d}=\mathbb{E}_{X\sim p_{\text{data}}}\left[\max(0,1-D(X))\right]+\mathbb{E}_{\hat{X}_{0}\sim p_{g}}\left[\max(0,1+D(\hat{X}_{0}))\right],$ (12)
where $p_{\text{data}}$ denotes the empirical distribution of real images from the training dataset, and $p_{g}$ represents the generator’s learned distribution, consistent with the notation introduced in Section 2. This loss encourages the discriminator to assign high scores to real images and low scores to generated images, stabilizing the adversarial training process.
Adversarial Training Objective. The adversarial loss is designed to encourage the generator to produce images that can maximize the discriminator’s output. Meanwhile, when the generator is trained with samples from larger timesteps, the predicted $x_{0}$ tends to be less accurate compared with predictions from smaller timesteps. To stabilize training and prevent the adversarial loss from dominating during these less reliable steps, we introduce a weighting mechanism. Specifically, we compute a scalar weighting adversarial signal as the square of the current timestep’s noise scale, i.e., $\omega(t)=\sigma_{t}^{2}$, and scale the adversarial loss. Thus, the adversarial loss for the generator is:
$\mathcal{L}_{g}=-\omega(t)\cdot\mathbb{E}_{\hat{X}_{0}\sim p_{g}}\left[D(\hat{X}_{0})\right]=-\sigma_{t}^{2}\cdot\mathbb{E}_{\hat{X}_{0}\sim p_{g}}\left[D(\hat{X})\right].$ (13)
This design ensures that the generator focuses more on the DMD gradient during noisy, high-timestep stages—where adversarial feedback may be unreliable—and benefits more from GAN guidance at cleaner, low-noise steps. In practice, this improves training stability and overall sample quality.
4 Experimental Results
4.1 Experimental Setup
Datasets. Following DMD2 yin2024improved , our experiments are conducted using a filtered set of the LAION-5B schuhmann2022laion dataset, which provides high-quality image-text pairs for training. We select images with a minimum aesthetic score (aes score) of 5.0 and a shorter dimension of at least 1024 pixels, ensuring the dataset comprises visually appealing, high-resolution images suitable for our model’s requirements.
For evaluation, we construct a validation set using the COCO 2017 lin2014coco validation set, which contains 5,000 images. Each image in this set is paired with the text annotation that yields the highest CLIP Score (ViT-B/32), thus forming a robust text-image validation set. We also evaluate compositional generation using T2I-CompBench huang2023t2i , a benchmark spanning attribute binding, object relationships, and complex compositions, which is designed to test models on generating semantically coherent images with diverse object interactions.
Text-to-Image Models. We conduct extensive experiments on three representative large-scale text-to-image models: FLUX.1 dev (12B) flux2024 , Stable Diffusion 3.5 Large (8B) esser2024scaling , and SDXL (2.6B) podell2023sdxl , which span different model sizes and generative paradigms. Results demonstrate the generality and effectiveness of our method across both flow-based and conventional diffusion architectures.
Evaluation Metrics. Following  wang2024phased ; lin2024sdxl ; yin2024improved , we report FID and Patch FID of all baselines and the generated images of original teacher models to assess distillation performance and high-resolution details, dubbed FID-T and Patch FID-T. We also report CLIP Score (ViT-B/32) to evulate text-image alignment and further include some recently proposed metrics, such as HPS v2 wu2023human , ImageReward imagereward , and PickScore pickscore to offer a more comprehensive evaluation of the model performance.
4.2 Text to Image Generation
Comparison Baselines. For the distillation of SDXL, we compare our method with baselines including LCM luo2023latent , PCM wang2024phased , Flash Diffusion chadebec2025flash , SDXL-Lightning lin2024sdxl , Hyper-SD ren2024hyper , and DMD2 yin2024improved . As for SD 3.5 Large, we compare our method with SD 3.5 Large Turbo sauer2024adversarial . For FLUX.1 dev, we compare with Hyper-FLUX ren2024hyper , FLUX.1 schnell flux2024 , and FLUX-Turbo-Alpha alimama2024flux1turboalpha .
Table 1: Quantitative Results on COCO-5K Dataset. Bold: best. Underline: second best. Our proposed approaches superior distillation performance accross different models on 4-step generation.
Method # NFE $\downarrow$ FID-T $\downarrow$ Patch FID-T $\downarrow$ CLIP $\uparrow$ HPSv2 $\uparrow$ Pick $\uparrow$ ImageReward $\uparrow$
Stable Diffusion XL Comparison
SDXL podell2023sdxl 80 0.3293 0.2930 22.67 0.8719
LCM-SDXL luo2023latent 4 18.47 30.63 0.3230 0.2824 22.22 0.5693
PCM-SDXL wang2024phased 4 14.38 17.77 0.3242 0.2920 22.54 0.6926
Flash-SDXL chadebec2025flash 4 17.97 23.24 0.3216 0.2830 22.17 0.4295
SDXL-Lightning lin2024sdxl 4 13.67 16.57 0.3214 0.2931 22.80 0.7799
Hyper-SDXL ren2024hyper 4 13.71 17.49 0.3254 0.3000 22.98 0.9777
DMD2-SDXL yin2024improved 4 15.04 18.72 0.3277 0.2963 22.98 0.9324
Ours-SDXL 4 17.76 21.01 0.3248 0.3010 23.17 0.9951
Stable Diffusion 3.5 Comparison
SD 3.5 Large esser2024scaling 100 0.3310 0.2993 22.98 1.1629
SD 3.5 Large Turbo sauer2024adversarial 4 13.58 22.88 0.3262 0.2909 22.89 1.0116
Ours-SD 3.5 4 13.38 17.48 0.3286 0.3016 23.01 1.1713
Ours-SD 3.5 (Euler) 4 15.24 20.26 0.3287 0.3008 22.90 1.2062
FLUX Comparison
FLUX.1 dev flux2024 50 0.3202 0.3000 23.18 1.1170
25 0.3207 0.2986 23.14 1.1063
FLUX.1-schnell flux2024 4 0.3264 0.2962 22.77 1.0755
Hyper-FLUX ren2024hyper 4 11.24 23.47 0.3238 0.2963 23.09 1.0983
FLUX-Turbo-Alpha alimama2024flux1turboalpha 4 11.22 24.52 0.3218 0.2907 22.89 1.0106
Ours-FLUX 4 15.64 19.60 0.3167 0.2997 23.13 1.0921
Ours-FLUX (Euler) 4 16.50 20.29 0.3171 0.3008 23.26 1.1424
Table 2: 4-Step Results on T2I-CompBench. Bold, Underline: best and second best in distilling the same teacher. Our distilled SD 3.5 model approaches state-of-the-art distillation performance.
Method Color Shape Texture Spatial Non-spatial Complex-3-in-1
LCM-SDXL luo2023latent 0.5997 0.4015 0.4958 0.1672 0.3010 0.3364
SDXL-Lightning lin2024sdxl 0.5758 0.4492 0.5154 0.2124 0.3098 0.3517
Hyper-SDXL ren2024hyper 0.6435 0.4732 0.5581 0.2213 0.3104 0.3301
PCM-SDXL wang2024phased 0.5591 0.4142 0.4693 0.2013 0.3099 0.3234
DMD2-SDXL yin2024improved 0.5811 0.4477 0.5175 0.2124 0.3098 0.3301
Ours-SDXL 0.6867 0.4828 0.5989 0.2224 0.3100 0.3594
SD 3.5 Large Turbo sauer2024adversarial 0.7050 0.5443 0.6512 0.2839 0.3130 0.3520
Ours-SD 3.5 0.7657 0.6069 0.7427 0.2970 0.3177 0.3916
Ours-SD 3.5 (Euler) 0.7711 0.6149 0.7543 0.2857 0.3182 0.3968
FLUX.1 schnell flux2024 0.7317 0.5649 0.6919 0.2626 0.3122 0.3669
Hyper-FLUX ren2024hyper 0.7465 0.5023 0.6153 0.2945 0.3116 0.3766
FLUX-Turbo-Alphaalimama2024flux1turboalpha 0.7406 0.4873 0.6024 0.2501 0.3094 0.3688
Ours-FLUX 0.7284 0.5055 0.6031 0.2451 0.3028 0.3652
Ours-FLUX (Euler) 0.7363 0.5120 0.6112 0.2521 0.3028 0.3697
Quantitative Comparison. The 4-step comparison results on COCO-5K and T2I-CompBench are presented in Tab. 1 and Tab. 2, respectively. For flow-matching models, we report both stochastic and deterministic sampling results, denoted as “Ours” and “Ours (Euler)”. As shown in Tab. 1, our method consistently outperforms previous distillation baselines across a wide range of metrics. On SD 3.5, both “Ours-SD 3.5” and “Ours-SD 3.5 (Euler)” achieve the best and second-best scores on all metrics, even surpassing the teacher model in HPSv2, PickScore, and ImageReward. On SDXL, our method ranks first in HPSv2, PickScore, and ImageReward, with a marginal drop in text-image alignment. For FLUX.1 dev, our models again deliver top performance across several metrics. The strong results under both stochastic and deterministic settings also confirm the robustness of our approach. In terms of T2I-CompBench, the results in Tab. 2 demonstrate that “Ours-SD 3.5 (Euler)” achieves state-of-the-art performance across all evaluated methods in five dimensions—color, shape, texture, spatial, non-spatial consistency, and the “Complex-3-in-1” metric. These results highlight the fine-grained fidelity and superior attribute alignment of our approach. “Ours-SDXL” also achieves the best performance in five out of the six evaluated metrics for SDXL distillation, the highest among compared methods. Further results and detailed analyses are provided in the appendix.
Qualitative Comparison. Fig 5 presents qualitative comparisons across a set of prompts. Our method generates images with sharper details, better limb structure, and more coherent lighting dynamics, compared to teacher models and baselines. Notably, “Ours-SD3.5" and “Ours-FLUX" produce more faithful and photorealistic generations under challenging prompts involving fine textures, human faces, and scene lighting. Fig. 6 also presents examples of our method on SD 3.5 Large. Additional qualitative results and discussion are provided in the appendix.
Figure 5: Qualitative comparisons on challenging prompts across methods. Our method shows superior fidelity, especially in rendering human faces, scene composition, and fine-grained textures.
4.3 Ablation Studies
Table 3: Ablation Study Results of IDA, ISG, and VFM Discriminator.
[…]
Stable Diffusion 3.5 Large
Ours 13.58 0.3288 0.2989 5.4559
wo ISG 17.00 0.3246 0.2971 5.4527
wo ISG, wo IDA 43.04 0.3000 0.2555 5.1018
Stable Diffusion XL
DMD2-SDXL yin2024improved 15.04 0.3277 0.2964 5.5305
DMD2 w VFM 18.55 0.3234 0.2998 5.6252
[…]
figureThe ISG improves training consistency, especially in the early stage of training.
Figure 6: 1024×1024 samples produced by our 4-step generator distilled from SD 3.5 Large.
Effectiveness of Implicit Distribution Alignment. To assess the effectiveness of our proposed IDA strategy, we conduct experiments on SD 3.5 Large with various TTUR ratios. As shown in Fig. 3, we compare FID curves across different settings, both with and without IDA. Without IDA, the curves corresponding to “TTUR(5)”, “TTUR(10)”, and “TTUR(20)” exhibit severe oscillations, indicating unstable training dynamics and unreliable optimization of the fake distribution—even at a high ratio of 20:1. This instability leads to inaccurate DMD gradients and poor convergence. In contrast, the settings that incorporate IDA (i.e., “IDA+TTUR(5)” and “IDA+TTUR(10)”) demonstrate significantly smoother and more stable FID reductions, highlighting IDA’s ability to stabilize training and improve convergence, even at a relatively small TTUR ratio (5:1).
In addition to the FID analysis, we report quantitative comparisons in Tab. 3 between “w/o ISG” and “w/o ISG, w/o IDA” using four metrics: FID-T, CLIP Score, HPSv2, and AESv2. Across all metrics, adding IDA leads to consistent improvements, further confirming that IDA plays a key role in enhancing training stability and distillation quality.
Intra-Segment Guidance. To evaluate the effectiveness of the Intra-Segment Guidance (ISG) module during distillation, we conduct an ablation study on Stable Diffusion 3.5 Large. As shown in Tab. 3, we compare our model with and without ISG (denoted as “Ours” and “w/o ISG”, respectively) on the COCO-5K dataset. The results indicate that incorporating ISG leads to significant improvements across all aspects, including image quality, text-image alignment, and human preference quality.
In addition, Fig. 3 presents a qualitative comparison at 3K training iterations, during which the generators have been updated for only 300 steps under 10:1 TTUR ratio. We observe that the model trained with ISG produces visually more consistent and semantically accurate images even at early training stages, whereas the model without ISG suffers from noticeable color shifts and degraded image fidelity. This highlights ISG’s contribution to training stability and convergence efficiency.
VFM-Based Discriminator. To assess the benefit of integrating Vision Foundation Model (VFM)-based discriminator, we conduct comparative experiments on the SDXL backbone. As shown in Tab. 3, we compare the DMD2-SDXL—equipped with a diffusion-based discriminator—with our method using the VFM discriminator (denoted as “DMD2 w VFM”). Across multiple evaluation metrics, “DMD2 w VFM” achieves better human preference alignment and aesthetic quality. These results demonstrate that the VFM-based discriminator provides stronger visual priors to the generator.
5 Related Work
Diffusion Distillation methods mainly fall into two categories: trajectory-based and distribution-based approaches. Trajectory-based methods, such as Direct Distillation luhman2021knowledge and Progressive Distillation salimans2022progressive ; ren2024hyper ; lin2024sdxl ; chadebec2025flash , learn to replicate the denoising trajectory, while Consistency Models song2023consistency ; luo2023latent ; kimctm ; wang2024phased ; lu2024simplifying ; chen2025sana enforce self-consistency across steps. Distribution-based methods aim to match the generative distribution, including GAN-based distillation Wang2023DiffusionGAN ; wang2024efficient ; luo2024you and VSD variants wang2023prolificdreamer ; yin2024one ; yin2024improved . ADD sauer2024adversarial and LADD sauer2024fast explored distilling diffusion models using adversarial training with pretrained feature extractors. RayFlow shao2025rayflow explored sampling important timesteps for better distillation. Among these, DMD2 yin2024improved has shown strong results on standard diffusion models (e.g., SDXL), but its stability degrades on large-scale models. Our work builds upon DMD2 and addresses these limitations by introducing SenseFlow, which scales distribution matching distillation to SD 3.5 and FLUX.1 dev through improved alignment and regularization strategies.
6 Conclusions and Limitations
We scale up distribution matching distillation for large flow-based models by introducing implicit distribution alignment and intra-segment guidance. Together with a VFM-based discriminator, these enhancements enable our model SenseFlow to achieve stable and effective few-step generation on both diffusion and flow-matching backbones. Our SD 3.5-based SenseFlow achieves state-of-the-art 4-step generation performance across all evaluated distillation methods, demonstrating its effectiveness on large-scale models. Meanwhile, its performance under more aggressive settings (e.g., 2-step, 1-step) and with alternative vision backbones oquab2023dinov2 ; sam2 ; amradio ; he2022mae remains unexplored. Finally, like other generative models, SenseFlow raises concerns regarding potential misuse and labor displacement, underscoring the importance of responsible deployment.
Appendix A Appendix
A.1 Broader Impact
Our work focuses on improving the efficiency and quality of text-to-image diffusion models, particularly on large-scale architectures. This has several potential societal impacts, both positive and negative. On the positive side, the proposed distillation framework significantly accelerates the sampling process of large models such as FLUX.1 dev and SD 3.5 Large, making high-quality image synthesis more accessible and practical for real-world applications. These improvements can benefit a wide range of domains, including education, digital content creation, scientific visualization, and assistive design tools, by enabling faster, more cost-efficient generation of customized visual content.
However, similar to other text-to-image models, our method inherits risks associated with generative models. These include the potential misuse of fast image synthesis for generating fake content, spreading misinformation, or fabricating identities. Additionally, like many generative models, our distilled networks are susceptible to reflecting biases present in the training data, which may result in unfair or unrepresentative outputs. As a future direction, we are interested in investigating methods for detecting and mitigating such biases in diffusion models, building on recent work in fairness-aware generation. We also plan to introduce clear usage guidelines and responsible deployment practices, including detailed user manuals, to promote ethical and transparent use of the technology.
A.2 Implementation Details
Our entire framework is implemented in PyTorch with CUDA acceleration and is trained using 8 A100 GPUs with a total batch size of 8. We adopt the AdamW optimizer loshchilov2017decoupled with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate is set to $61e-$ 6 for the distillation of SDXL and SD 3.5 Large, and $51e-$ 5 for FLUX.1 dev. To efficiently support large-scale model training, we utilize Fully Sharded Data Parallel (FSDP), which enables memory-efficient and scalable training of large models.
Figure 7: Design of the VFM-based discriminator.
Timestep settings.
We adopt different coarse timestep schedules depending on the model architecture. For SDXL, we follow the 1000-step discrete DDPM schedule used in DMD2 yin2024improved , selecting step indices $\{249,499,749,999\}$. For SD 3.5 Large, we switch to continuous timestep values $\{0.25,0.5,0.75,1.0\}$ , which are more suitable for flow-based models. In the case of FLUX.1 dev, which adopts a shifted $\sigma$ inference strategy, we directly use the corresponding sigmas $\{0.512,0.759,0.904,1.0\}$ as coarse anchors.
Training details.
We set the default TTUR (Two Time-Scale Update Rule) ratio to $5$ in our main experiments on SDXL, SD 3.5 Large, and FLUX.1 dev. For large flow-based models such as SD 3.5 Large and FLUX.1 dev, we apply all proposed improvements, including Implicit Distribution Alignment (IDA), Intra-Segment Guidance (ISG), and the VFM-based Discriminator. For the diffusion-based SDXL model, we employ ISG and the VFM-based Discriminator while omitting IDA.
A.3 Detailed VFM-Based Discriminator Design
As shown in Fig. 7, the discriminator integrates pretrained vision (DINOv2) and language (CLIP) encoders to provide semantically rich and spatially aligned supervision. Given an input image $x$, we apply normalization (from $[-1,1]$ to $[0,1]$) and differentiable data augmentation (including color jitter, translation, and cutout). The augmented image is processed by a frozen DINOv2 vision transformer to extract multi-level semantic features. Each selected layer output is reshaped into a 2D spatial map (e.g., $[B,C,H,W]$) and passed through a lightweight convolutional head composed of spectral-normalized residual blocks.
A reference image $x_{\text{ref}}$ is processed through the same DINOv2 pathway (without augmentation) to extract corresponding semantic features. Meanwhile, the text prompt is encoded by a CLIP (ViT-L/14) text encoder into a condition feature $c$, which is projected to a spatial map. Each discriminator head fuses the image feature, reference feature, and prompt condition via element-wise multiplication and spatial summation to compute the final logits. (Note: In Section 3.4, we described the reference features $r$ as extracted by the CLIP encoder. In practice, $r=f_{\mathrm{VFM}}(x_{\text{ref}})$ is obtained using the same DINOv2 backbone as the input image. The Fig. 2 should also be corrected.)
A.4 Training Algorithm
To more clearly illustrate our training process, we provide the full algorithmic details in Algorithm 1. We adopt model-specific hyperparameter settings for better distillation performance. In particular, we set the hyperparameter $\lambda_{\text{IDA}}$ of implicit distribution alignment to $0.97$ by default. For the intra-segment guidance loss, $\lambda_{\text{ISG}}$ is set to $0.2$ for SDXL, and $1.0$ for both SD 3.5 and FLUX.1 dev.
Algorithm 1 SenseFlow Training Algorithm
1:pretrained teacher model $\mu_{\text{real}}$, real dataset $\mathcal{D}_{\text{real}}$, generator update frequency $f$, coarse timestep set $S=\{\tau_{0},\tau_{1},\tau_{2},\tau_{3}\}$
2:trained few-step generator $G$
3:$G\leftarrow\text{copyWeights}(\mu_{\text{real}})$triangleright Initialize generator
4:$\mu_{\text{fake}}\leftarrow\text{copyWeights}(\mu_{\text{real}})$triangleright Initialize fake distribution network
[…]
6:for iteration $=1$ to max_iters do
7:     $z\sim\mathcal{N}(0,I)$
8:     Sample $\tau_{i}$ from $S$triangleright Pick timestep for current iteration
9:     Sample $x_{\text{real}}\sim\mathcal{D}_{\text{real}}$
10:     if $random()<0.5$ then $\triangleright$ With 50% probability, use backward simulation
11:         $x_{\tau_{i}}\leftarrow\text{multiStepSampling}(z,\tau_{3}\to\tau_{i}))$
12:     else
13:         $x_{\tau_{i}}\leftarrow\text{forwardDiffusion}(x_{\text{real}},\tau_{i})$
14:     end if
15:     $x\leftarrow G(x_{\tau_{i}})$
16:     if iteration mod $f=0$ then
17:         $\mathcal{L}_{\text{DMD}}\leftarrow\text{distributionMatching}(\mu_{\text{real}},\mu_{\text{fake}},x)$
18:         $\mathcal{L}_{\text{G}}\leftarrow-\sigma_{\tau_{i}}^{2}\cdot\mathbb{E}[D(x)]$triangleright Eq. 13 $\triangleright$ Intra-segment guidance (ISG)
19:         $t_{\text{mid}}\sim\mathcal{U}(\tau_{i},\tau_{i-1})$
20:         $x_{\text{mid}}\leftarrow\mu_{\text{real}}(x_{\tau_{i}},\tau_{i}\to t_{\text{mid}})$
21:         $x_{\text{tar}}\leftarrow G(x_{\text{mid}},t_{\text{mid}}\to\tau_{i-1})$
22:         $x_{\tau_{i-1}}\leftarrow G(x_{\tau_{i}},\tau_{i}\to\tau_{i-1})$
23:         $\mathcal{L}_{\text{ISG}}\leftarrow\text{MSE}(x_{\tau_{i-1}},\text{stopgrad}(x_{\text{tar}}))$
24:         $\mathcal{L}_{\text{G}}\leftarrow\mathcal{L}_{\text{DMD}}+\lambda_{\text{G}}\cdot\mathcal{L}_{\text{G}}+\lambda_{\text{ISG}}\cdot\mathcal{L}_{\text{ISG}}$triangleright Final loss function for generator
25:         $G\leftarrow\text{update}(G,\mathcal{L}_{\text{G}})$triangleright Implicit distribution alignment (IDA), as in Eq. 8
26:         $\mu_{fake}\leftarrow\text{IDA}(G,\mu_{fake},\lambda_{\text{IDA}})$
27:     end $\triangleright$ Update fake score network $\mu_{\text{fake}}$
[…]
29:     $x_{t}\leftarrow\text{forwardDiffusion}(\text{stopgrad}(x),t)$
30:     $\mathcal{L}_{\text{denoise}}\leftarrow\text{denoisingLoss}(\mu_{\text{fake}}(x_{t},t),\text{stopgrad}(x))$
31:     $\mu_{\text{fake}}\leftarrow\text{update}(\mu_{\text{fake}},\mathcal{L}_{\text{denoise}})$triangleright Update discriminator $D$
32:     $\mathcal{L}_{\text{D}}\leftarrow\mathbb{E}[\max(0,1-D(x_{\text{real}})]+\mathbb{E}[\max(0,1+D(x)]$triangleright Eq. 12
33:     $D\leftarrow\text{update}(D,\mathcal{L}_{\text{D}})$
34:end for
A.5 More Experimental Results
Effect of Different Adversarial Loss Weights.
In our main experiments, the hyperparameter $\lambda_{\text{G}}$ in Algorithm 1, Line 22, is set to 0.5, 0.1, and 2.0 for SDXL, SD 3.5 Large, and FLUX.1 dev, respectively. To further investigate the impact of this hyperparameter, we conduct an ablation study using SDXL as an example, decreasing $\lambda_{\text{G}}$ to 0.25. The results are presented in Tab.4. We observe that setting $\lambda_{\text{G}}=0.5$ leads to improved performance across most metrics, including CLIP Score, HPSv2, PickScore, and ImageReward. Notably, this configuration achieves the best scores on HPSv2, PickScore, and ImageReward among all methods in Tab.1. These results highlight the strong semantic and visual supervision capabilities of our VFM-based discriminator.
Table 4: Quantitative Results of different backbone scales.
Method FID-T $\downarrow$ CLIP Score $\uparrow$ HPSv2 $\uparrow$ Pick Score $\uparrow$ ImageReward $\uparrow$
Hyper-SDXL ren2024hyper 13.71 0.3254 0.3000 22.98 0.9777
Ours ($\lambda_{\text{G}}=0.25$) 17.53 0.3234 0.3003 23.15 0.9326
Ours ($\lambda_{\text{G}}=0.5$) 17.76 0.3248 0.3010 23.17 0.9951
Results of Different Backbone Scales.
We evaluate the impact of different VFM backbone scales (ViT-S, B, and L) in the discriminator on SDXL distillation. Interestingly, the results (Tab.1) do not follow a monotonic trend with respect to model size. ViT-B achieves the best FID-T, while ViT-S yields higher CLIP Score and ImageReward. ViT-L slightly outperforms others on HPSv2 and PickScore. These findings suggest that different backbone scales offer different trade-offs in semantic alignment versus visual fidelity, and that larger backbones do not necessarily guarantee consistent improvements across all metrics. This observation is partially consistent with findings in the ADDsauer2024adversarial paper, which also noted diminishing returns when scaling the discriminator. In our main paper, we adopt ViT-L as the default backbone for the VFM-based discriminator.
Table 5: Quantitative Results of different backbone scales.
Method FID-T $\downarrow$ CLIP Score $\uparrow$ HPSv2 $\uparrow$ Pick Score $\uparrow$ ImageReward $\uparrow$
Ours w ViT-S 17.26 0.3262 0.2983 23.12 0.9635
Ours w ViT-B 16.58 0.3234 0.2991 23.07 0.9218
Ours w ViT-L 17.53 0.3239 0.3003 23.15 0.9326
Examples from T2I-CompBench.
As shown in Fig. 8, we present visual comparisons of different methods on SDXL using the T2I-CompBench benchmark. These qualitative results clearly highlight the superiority of our approach across multiple aspects, including color fidelity (rows 1 and 2), shape consistency (row 3), material and texture (row 4), and complex spatial arrangements (row 5). Additionally, we also present more examples of our method on SDXL in Fig. 9.
Figure 8: Examples from T2I-CompBench.
Figure 9: 1024×1024 samples produced by our 4-step generator distilled from SDXL.
A.6 Prompts for Fig. 1, Fig. 6, and Fig. 9
We use the following prompts for Fig. 1. From left to right, top to bottom:
  • A red fox standing alert in a snow-covered pine forest
  • A girl with a hairband performing a song with her guitar on a warm evening at a local market, children’s story book
  • Astronaut on a camel on mars
  • A cat sleeping on a windowsill with white curtains fluttering in the breeze
  • A stylized digital art poster with the word "SenseFlow" written in flowing smoke from a stage spotlight
  • A surreal landscape inspired by The Dark Side of the Moon, with floating clocks and rainbow beams
  • a hot air balloon in shape of a heart. Grand Canyon
  • A young man with a leather jacket and messy hair playing a cherry-red electric guitar on a rooftop at sunset
  • A young woman wearing a denim jacket and headphones, walking past a graffiti wall
  • A photographer holding a camera, squatting by a lake, capturing the reflection of the mountains in an early morning
  • a young girl playing piano
  • A close-up of a woman’s face, lit by the soft glow of a neon sign in a dimly lit, retro diner, hinting at a narrative of longing and nostalgia
Besides, we use the following prompts for Fig. 6. From left to right, top to bottom:
  • A quiet room with Oasis album covers framed on the wall, acoustic guitar resting on a stool
  • An astronaut lying in the middle of white ROSES, in the style of Unsplash photography.
  • cartoon dog sits at a table, coffee mug on hand, as a room goes up in flames. "Help" the dog is yelling
  • Art illustration, sports minimalism style, fuzzy form, black cat and white cat, solid color background, close-up, pure flat illustration, extreme high-definition picture, cat’s eyes depict clear and meticulous, high aesthetic feeling, graphic, fuzzy, felt, minimalism, blank space, artistic conception, advanced, masterpiece, minimalism, fuzzy fur texture.
  • Close-up of the top peak of Aconcagua, a snow-covered mountain in the Himalayas at sunrise during the golden hour. Award-winning photography, shot on a Canon EOS R5 in the style of Ansel Adams.
  • A curvy timber house near a sea, designed by Zaha Hadid, represents the image of a cold, modern architecture, at night, white lighting, highly detailed
  • a teddy bear on a skateboard in times square
  • a black and white picture of a woman looking through the window, in the style of Duffy Sheridan, Anna Razumovskaya, smooth and shiny, wavy, Patrick Demarchelier, album covers, lush and detailed
As for Fig. 9, we use following prompts from left to right, top to bottom:
  • Astronaut in a jungle, cold color palette, muted colors, detailed, 8k
  • A bookshelf filled with colorful books, a potted plant, and a small table lamp
  • A dreamy beachside bar at dusk serving mojitos and old fashioneds, guitars hanging on the wall
  • A portrait of a human growing colorful flowers from her hair. Hyperrealistic oil painting. Intricate details.
  • Peach-faced lovebird with a slick pompadour.
  • a stunning and luxurious bedroom carved into a rocky mountainside seamlessly blending nature with modern design with a plush earth-toned bed textured stone walls circular fireplace massive uniquely shaped window framing snow-capped mountains dense forests
  • An acoustic jam session in a small café, handwritten setlist on the wall, cocktails on every table
  • a blue Porsche 356 parked in front of a yellow brick wall.
A.7 Licenses for existing assets
We use only publicly available and properly licensed open-source datasets and pretrained models in this work. All assets are cited in the main paper, and their licenses explicitly permit academic usage, redistribution, or derivative works under specific conditions. Below is a list of the key assets used and their associated licenses:
  • LAION-5B: Licensed under CC-BY 4.0.
    A large-scale text-image dataset used in pretraining and evaluation contexts.
  • COCO-2017: Licensed under a custom non-commercial research license.
    Commonly used for generation evaluation.
  • Stable Diffusion XL: Licensed under CreativeML Open RAIL++-M.
    Used as a diffusion based teacher model in our distillation framework.
  • Stable Diffusion 3.5: Licensed under CreativeML Open RAIL++-M.
    Used as a large flow-matching base model.
  • FLUX.1-dev: Licensed under CreativeML Open RAIL++-M.
    Used as a large flow-matching base model.
  • DINOv2: Licensed under Apache 2.0.
    Used as the frozen vision foundation backbone in our discriminator design.
  • OpenCLIP: Licensed under Apache 2.0.
    Serves as the text encoder for prompt conditioning in the discriminator.
  • T2I-CompBench: Licensed under the MIT License.
    Used for benchmark comparison of compositional generation performance.
All assets were used in accordance with their respective licenses.